Problem 3. Let (Xi, X2, X3, X4) be Multinomial(n, 4,1/6, 1/3,1/8,3/8). Derive the joint mass function of the pair (X3, X4). You should be able to do this with almost no computation.

I will rate asap. Thanks
Answer #1

We are given here that:
P(X3 = 1) = 1/8 and P(X4 = 1) = 3/8

Therefore, P(X3 = X4 = 0) = 1 - (1/8) - (3/8) = 1/2

Therefore, the joint mass function of X3, X4 here is obtained as:

P(X3 = x3, X4 = x4) = Probability of getting x3 instances of X3, x4 instances of X4 and (4 - x3 - x4) instances of anything other than x3 / x4.

Therefore, the probability now is obtained here as:

P(X_3= x_3, X_4 = x_4) = \binom{4}{x_3}(1/8)^{x_3}\binom{4 - x_3}{x_4}(3/8)^{x_4}*(1/2)^{4 - x_3 - x_4}

This is the required joint PDF for X3 , X4